# Neural Uncertainty Principle: A Unified View of Adversarial Fragility and LLM Hallucination

### 20 Mar 2026 v1

@@ I Introduction
Here, we take a unified view and argue that both failure modes share a geometric signature: the model enters an boundary-stress of the loss landscape. In adversarial attacks, small input perturbations can traverse locally sharp loss contours; in LLMs, weak prompt conditioning leaves a large set of continuations similarly compatible before decoding, enabling prior-driven drift. We show that these are two opposite ways of mismanaging the same “uncertainty budget”—an imbalance between input-space localization and gradient-space sensitivity.
@@ III-A NUP in a Nutshell: Roadmap From Geometry to Experiments
  • Exp. 5–6 (LLM intervention): low $c_{\text{prompt}}$ indicates under-conditioning at prefill; selecting higher-$c_{\text{prompt}}$ prompt variants moves behavior away from the lower-right corner toward the intermediate band.
@@ III-B Loss-weighted state and canonical directional operators
so that $|\psi_{c}(x)|^{2}=A_{c}(x)^{2}=w_{c}(x)$. We fix the phase scale $\alpha=1$ throughout. It only sets the phase convention of the auxiliary state and is not treated as a tunable parameter. For a (densely defined) operator $\hat{O}$ on $L^{2}(\mathcal{X})$, we write its expectation as [^2] The expectation $\langle\hat{O}\rangle_{c}$ is the average value of this quantity in the state $\psi_{c}$, computed via the usual $L^{2}$ inner product.
@@ IV Experimental Protocol
The preceding section derived four testable predictions from the NUP. We design Experiments 1–6 to falsify these predictions across two modalities: discriminative vision (classification) and generative language (reasoning). Detailed training recipes and hyperparameters are deferred to Supplementary Material.
@@ IV-A Datasets and Model Zoo
  • Transformers: Vision Transformer ViT-Tiny/Small, Swin-Transformer.
  • State-Space Models: Vision Mamba [72] Vim-Tiny (Mamba-based vision backbone).
@@ IV-B Operationalizing the Conjugate Probe
For an image input $X\in\mathbb{R}^{C\times H\times W}$, we compute the gradient of the task loss $\mathcal{L}$ with respect to the input, $P=\nabla_{X}\mathcal{L}$. Both $X$ and $P$ are flattened into 1-D vectors $x$ and $p$. Since standard vision pipelines normalize inputs (zero-mean), the sample-level conjugate correlation and its dataset-level average over $N$ samples are defined as:
@@
Remark. The NUP is expressed via an operator correlation $\rho_{c}(u)$ under a loss-induced state, while we use the practical proxy $\bar{c}_{\text{img}}$. $c_{\text{img}}$ is intended to track the input–gradient coupling channel that enters the Robertson–Schrödinger covariance term: under the loss-phase construction, $\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})$ reduces exactly to a (loss-weighted) scalar covariance between $x_{u}$ and $p_{u}$ (Theorem III.2), and the dot-product $x^{\top}p$ is the direction-averaged form of this coupling.
@@
For language models, we extract the signal during the prompt processing stage (pre-fill). We compute a mean-centered cosine between the prompt embeddings and their input gradients. Let $X\in\mathbb{R}^{T\times d}$ be the embedding tensor for the $T$ valid tokens, and $P\in\mathbb{R}^{T\times d}$ be the corresponding gradient tensor. We use a shifted loss, i.e., the next-token negative log-likelihood. Unless otherwise stated, the shifted NLL loss is computed on the non-system prefill tokens of the chat template (i.e., the user prompt together with template tokens available before answer generation), so the score is available at prefill time without generating any answer tokens; it still requires one backward pass with respect to the prompt embeddings. We flatten these tensors into 1-D vectors $x,p\in\mathbb{R}^{Td}$ and define the mean-centered vectors
@@ IV-C Evaluation Metrics
We treat hallucination detection as a binary classification task (($1=$ Hallucination, $0=$ Normal)), quantifying performance via the Area Under the ROC Curve (AUROC) with bootstrap resampling ($N_{\text{boot}}=2000$). For prefill prompt selection, we report the Top-1 Hit Rate (frequency of selecting the judge-preferred prompt) and Mean Regret (the average quality gap between the selected and the optimal answer). Deterministic decoding is used for reproducibility.
@@ V-A Experimental Roadmap
Our experiments falsify the NUP predictions using a simple, computable proxy—the CC-Probe—across vision classification and LLM prompting. We organize Exp. 1–6 into three phases [^3] To reproduce Exp. 1–6, please refer to the released artifacts at https://doi.org/10.57760/sciencedb.29613.:
@@
Figure 3: Effect of gradient-aligned perturbations on accuracy and $\bar{c}_{\text{img}}$. We evaluate ResNet-18, ViT, and Vim on CIFAR-10 and ImageNet-100. Left column: accuracy; right column: $\bar{c}_{\text{img}}$. We compare Clean, +FGSM and –FGSM across perturbation levels $\epsilon$ (see Supplementary Material for exact values).
@@
ResNet-18 Base 91.21 0.62 0.54 0.14
TRADES-PGD 74.57 44.14 43.92 42.27 100
ConjMask (Exp. 3) 89.50 83.96 83.35 0.07 20
ConjMask+LogitReg (Exp. 4) 83.62 67.69 55.77 41.81 30
ViT Base 83.37 0.00 0.01 0.00
TRADES-PGD 71.95 32.95 32.46 35.51 120
ConjMask (Exp. 3) 83.01 18.56 10.14 0.00 50
ConjMask+LogitReg (Exp. 4) 77.86 58.12 27.64 22.86 70
EfficientNet-B0 Base 88.86 3.76 3.09 0.03
TRADES-PGD 56.88 30.65 30.48 28.05 260
ConjMask (Exp. 3) 86.67 73.70 72.83 0.27 70
ConjMask+LogitReg (Exp. 4) 84.91 44.29 11.33 11.61 55
@@
We start from a standard checkpoint (trained as in Exp. 1) and use the released mask_pelnaty_ablation branch. For each mini-batch, we first compute a forward pass and mark samples as triggered if either the current prediction is incorrect, or the prediction confidence is below a fixed threshold (default $0.2$). For each triggered image $X$ (flattened as $x$) we compute the input gradient $P=\nabla_{X}\mathcal{L}$ (flattened as $p$). In the released implementation, this probing step intentionally switches the model to eval mode; moreover, this mode is kept for the subsequent forward passes of the same optimization step. Since the goal of ConjMask is to reduce the CC-Probe $c_{\text{img}}=|\cos(x,p)|$ (Eq. (33)), we focus on the components that contribute most to this coupling statistic.
@@
We report robust accuracy under $L_{\infty}$ attacks at $\epsilon=8/255$, using PGD-20 and Auto-PGD variants (CE and DLR losses), together with standard clean accuracy in Table I:
@@
    CE robustness can be narrow and loss-dependent. The same models that are exceptionally strong on PGD/APGD-CE can collapse under APGD-DLR: ResNet-18 drops to $0.07\%$ and EfficientNet-B0 to $0.27\%$. This “strong on CE, weak on DLR” pattern highlights objective-specific geometry channels, which is consistent with the NUP view that manipulating one coupling channel may not cover all attack objectives.
@@
In this experiment we keep the same ConjMask masking procedure as Exp. 3 and add two lightweight auxiliary stabilizers on top of it. Concretely, the current code release activates a logit-side regularization branch (logits_loss) and an auxiliary-view consistency branch (random_loss), but their exact realizations differ across architectures (detailed in the Supplementary Material). We therefore use LogitReg as an umbrella term for these output-side stabilizers, rather than as a single architecture-invariant closed-form loss. The common goal is to complement the CE-gradient coupling channel targeted by ConjMask with score-space stabilization and auxiliary-view consistency.
We evaluate under the same $L_{\infty}$ setting as Exp. 3 ($\epsilon=8/255$) using PGD-20, APGD-CE, and APGD-DLR:
@@
    lightweight stabilizers restore DLR robustness with practical efficiency. Adding logit regularization and architecture-dependent auxiliary-view consistency on top of the same ConjMask procedure substantially improves APGD-DLR robustness where Exp. 3 collapses. On ResNet-18, APGD-DLR increases from $0.07\%$ (Exp. 3) to $41.81\%$ (Exp. 4), matching TRADES-PGD’s $42.27\%$, while maintaining high clean accuracy ($83.62\%$) and requiring no adversarial training, hence lower practical cost. The resulting robustness profile is more balanced than Exp. 3, with non-trivial performance retained under PGD-20 and APGD-CE ($67.69\%$ and $55.77\%$).
@@
    The gain is architecture-dependent, reflecting CE–DLR direction mismatch with partial overlap. For ViT and EfficientNet-B0, Exp. 4 trades part of the extreme PGD/APGD-CE robustness of Exp. 3 for improved APGD-DLR robustness (ViT: $0.00\%\to 22.86\%$; EfficientNet-B0: $0.27\%\to 11.61\%$), but the improvement is less pronounced than on ResNet-18 and comes with a noticeable drop in PGD-20/APGD-CE.
Overall, despite the architecture-dependent trade-off, our method remains practically usable since it achieves non-trivial robustness improvements with much less training time than AT.
@@
We falsify Proposition 4 by testing whether low prompt-level conjugate correlation signals hallucination risk in a trained LLM. For each of 500 math problems, we compute the prefill-stage probe on the prompt embeddings: we obtain the embedding tensor $x$ and its input gradient $p=\nabla_{x}\mathcal{L}(x)$ from a single backward pass, and form $c_{\text{prompt}}$ (Eq. (34)), which is computed before any answer tokens are generated.
@@
We report AUROC with bootstrap resampling ($N_{\text{boot}}=2000$) for distinguishing strict-consensus hallucinations ($P_{\text{vote}}=1.0$) from strict-consensus clean responses ($P_{\text{vote}}=0.0$), and also provide a relaxed-consensus analysis ($P_{\text{vote}}\geq 0.8$) as a robustness check. Results are summarized in Table II.
Two key findings emerge:
@@
TABLE III: (Exp. 6) Prefill-only prompt selection performance. All risk scores are computed before decoding any answer tokens. Regret quantifies the loss in answer quality when using the selected prompt compared to the oracle best prompt (lower is better).
Metric Selection Efficacy Best-vs-Rest Discrimination Spearman $\rho$
Top-1 Hit $\uparrow$ Mean Regret $\downarrow$ AUROC Score 95% CI vs. 0.5
Risk-Cos (Ours) 0.76 0.188 0.604 [0.55, 0.66] +20.8% -0.17
Risk-Margin 0.76 0.325 0.481 [0.43, 0.53] -3.8% -0.02
Risk-Entropy 0.57 0.657 0.381 [0.33, 0.43] -23.9% -0.18
Risk-Loss (NLL) 0.52 0.870 0.359 [0.31, 0.41] -28.2% -0.23
TABLE IV: (Exp. 6) Sensitivity analysis (Best-vs-Rest AUROC). Performance across three judge consensus scenarios.
Metric Hard Medium All
($P=1.0$) ($P\geq 0.6$) ($P\geq 0.2$)
Risk-Cos (Ours) 0.610 0.655 0.610
Risk-Margin 0.491 0.484 0.499
Risk-Entropy 0.516 0.446 0.468
Risk-Loss 0.426 0.381 0.422
@@
We use Perturbation-100, consisting of 100 problems with 5 semantically equivalent (paraphrased) prompt variants per problem. For each problem, we compute a prefill risk score for all five candidate prompts without decoding any answer tokens. We then select the single candidate with the lowest predicted risk and generate one answer from that selected prompt. For evaluation only, we also generate answers for all five prompt variants offline and score them with the same judge panel as in Exp. V-F. The highest-scoring variant defines the oracle best prompt for that problem.
@@
Each candidate prompt is assigned one scalar risk value (higher = worse) computed before decoding: Risk-Cos (ours), Risk-Margin, Risk-Entropy, and Risk-Loss (NLL) (definitions follow Sec. IV-B). All reported selection decisions in this experiment use only these prefill scores.
@@
We report three complementary metrics that separate top-choice accuracy from ranking stability:
  • Top-1 Hit Rate ($\uparrow$): the fraction of problems where the selected prompt equals the oracle best prompt.
  • Mean Regret ($\downarrow$): the average quality gap between the oracle best prompt and the selected prompt.
  • Best-vs-Rest AUROC ($\uparrow$): across all problems, treat each oracle best prompt as a positive instance and the remaining four variants as negatives; AUROC measures how well a risk score ranks the best prompt ahead of the rest (i.e., assigns it lower risk).
We also report Spearman correlation between risk-based rankings and judge-based rankings across the five variants per problem.

### 26 Mar 2026 v2

@@ I Introduction
Here, we take a unified view and argue that both failure modes share a geometric signature: the model enters a boundary-stress of the loss landscape. In adversarial attacks, small input perturbations can traverse locally sharp loss contours; in LLMs, weak prompt conditioning leaves a large set of continuations similarly compatible before decoding, enabling prior-driven drift. We show that these are two opposite ways of mismanaging the same “uncertainty budget”—an imbalance between input-space localization and gradient-space sensitivity.
@@ III-A NUP in a Nutshell: Roadmap From Geometry to Experiments
  • Exp. 5–6 (LLM intervention): low $c_{\text{prompt}}$ indicates under-conditioning at prefill; selecting higher-$c_{\text{prompt}}$ prompt variants moves behavior towards the Goldilocks Zone.
@@ III-B Loss-weighted state and canonical directional operators
so that $|\psi_{c}(x)|^{2}=A_{c}(x)^{2}=w_{c}(x)$. We fix the phase scale $\alpha=1$ throughout. It only sets the phase convention of the auxiliary state and is not treated as a tunable parameter. For a (densely defined) operator $\hat{O}$ on $L^{2}(\mathcal{X})$, we write its expectation [49] as  [^2] The expectation $\langle\hat{O}\rangle_{c}$ is the average value of this quantity in the state $\psi_{c}$, computed via the usual $L^{2}$ inner product.
@@ IV Experimental Protocol
The preceding section derived four testable predictions from the NUP. We design Experiments 1–6 to test these predictions across two modalities: discriminative vision (classification) and generative language (reasoning). Detailed training recipes and hyperparameters are deferred to Supplementary Material.
@@ IV-A Datasets and Model Zoo
  • Transformers: Vision Transformer ViT-Tiny and Swin-Tiny.
  • State-Space Models: Vision Mamba (Vim-Tiny) [73].
@@ IV-B Operationalizing the Conjugate Probe
For an image input $X\in\mathbb{R}^{C\times H\times W}$, we compute the gradient of the task loss $\mathcal{L}$ with respect to the input, $P=\nabla_{X}\mathcal{L}$. Both $X$ and $P$ are flattened into 1-D vectors $x$ and $p$. Because the practical vision probe is computed on the normalized input tensor fed to the model, we use the uncentered image-space cosine; the sample-level conjugate correlation and its dataset-level average over $N$ samples are defined as:
@@
Remark. The NUP is expressed via an operator correlation $\rho_{c}(u)$ under a loss-induced state, while we use the practical proxy $c_{\text{img}}$. $c_{\text{img}}$ is intended to track the input–gradient coupling channel that enters the Robertson–Schrödinger covariance term: under the loss-phase construction, $\mathrm{Cov}_{c}(\hat{x}_{u},\hat{p}_{u})$ reduces exactly to a (loss-weighted) scalar covariance between $x_{u}$ and $p_{u}$ (Theorem III.2), and the dot-product $x^{\top}p$ is the direction-averaged form of this coupling.
@@
For language models, we extract the signal during the prompt processing stage (prefill). We compute a mean-centered cosine between the prompt embeddings and their input gradients. Let $X\in\mathbb{R}^{T\times d}$ be the embedding tensor for the $T$ valid tokens, and $P\in\mathbb{R}^{T\times d}$ be the corresponding gradient tensor. We use a shifted loss, i.e., the next-token negative log-likelihood. Unless otherwise stated, the shifted NLL loss is computed on the non-system prefill tokens of the chat template (i.e., the user prompt together with template tokens available before answer generation), so the score is available at prefill time without generating any answer tokens; it still requires one backward pass with respect to the prompt embeddings. We flatten these tensors into 1-D vectors $x,p\in\mathbb{R}^{Td}$ and define the mean-centered vectors
@@ IV-C Evaluation Metrics
We treat Exp. 5 hallucination detection as a binary classification task (($1=$ hallucination, $0=$ non-hallucination)), quantified by the Area Under the ROC Curve (AUROC) with bootstrap resampling ($N_{\text{boot}}=2000$). For Exp. 6 prefill-stage prompt selection, the primary metrics are Top-1 Hit Rate (frequency with which the selected prompt belongs to the judge-best set) and Mean Regret (the average quality gap between the selected prompt and the best judge score within the same five-prompt cluster). Deterministic decoding is used for reproducibility.
@@ V-A Experimental Roadmap
Our experiments test the NUP predictions using a simple, computable proxy—the CC-Probe—across vision classification and LLM prompting. We organize Exp. 1–6 into three phases [^3] To reproduce Exp. 1–6, please refer to the released artifacts at https://doi.org/10.57760/sciencedb.29613.:
@@
Figure 3: Effect of gradient-aligned perturbations on accuracy and $\bar{c}_{\text{img}}$. We evaluate ResNet-18, ViT-Tiny, and Vim-Tiny on CIFAR-10 and ImageNet-100. Left column: accuracy; right column: $\bar{c}_{\text{img}}$. We compare Clean, +FGSM and –FGSM across perturbation levels $\epsilon$ (see Supplementary Material for exact values).
@@
ResNet-18 Base 91.52 0.01 0.02 0.10
TRADES-PGD 74.57 44.14 43.92 42.27 $\sim$ 100
ConjMask (Exp. 3) 85.27 77.55 77.02 1.28 $\sim$ 20
ConjMask+LogitReg (Exp. 4) 87.64 62.71 36.92 43.29 $\sim$ 30
ViT-Tiny Base 84.10 0.01 0.01 0.02
TRADES-PGD 71.95 32.95 32.46 35.51 $\sim$ 120
ConjMask (Exp. 3) 71.49 39.19 29.49 13.96 $\sim$ 45
ConjMask+LogitReg (Exp. 4) 71.24 43.15 14.59 15.91 $\sim$ 60
EfficientNet-B0 Base 88.14 0.24 0.17 0.00
TRADES-PGD 56.88 30.65 30.48 28.05 $\sim$ 260
ConjMask (Exp. 3) 83.91 69.10 68.28 0.20 $\sim$ 50
ConjMask+LogitReg (Exp. 4) 81.18 48.48 21.89 28.36 $\sim$ 65
@@
We start from a separately trained clean CIFAR-10 checkpoint for the same architecture. For each mini-batch, we first compute a forward pass and mark samples as triggered if either the current prediction is incorrect, or the prediction confidence is below a fixed threshold (default $0.2$). For each triggered image $X$ (flattened as $x$) we compute the input gradient $P=\nabla_{X}\mathcal{L}$ (flattened as $p$). In the released implementation, this probing step intentionally switches the model to eval mode; moreover, this mode is kept for the subsequent forward passes of the same optimization step. Since the goal of ConjMask is to reduce the CC-Probe $c_{\text{img}}=|\cos(x,p)|$ (Eq. (33)), we focus on the components that contribute most to this coupling statistic.
@@
We report robust accuracy under $L_{\infty}$ attacks at $\epsilon=8/255$, using PGD-20 and Auto-PGD variants (CE and DLR losses), together with standard clean accuracy in Table I. The following observations can be made:
@@
    CE robustness can remain objective-dependent. The same models that are strong on PGD/APGD-CE can still degrade sharply under APGD-DLR: for ConjMask (Exp. 3), APGD-DLR robust accuracy is $1.28\%$ on ResNet-18, $13.96\%$ on ViT-Tiny, and $0.20\%$ on EfficientNet-B0 (Table I). This “strong on CE, weaker on DLR” pattern highlights objective-specific geometry channels, which is consistent with the NUP view that manipulating one coupling channel may not cover all attack objectives.
@@
In Exp. 4, we keep the same ConjMask masking procedure as in Exp. 3 and add an output-side regularization term, denoted by LogitReg, together with an optional auxiliary-view consistency term. The purpose is to complement the CE-gradient coupling channel targeted by ConjMask with score-space stabilization, so that robustness is not restricted to CE-like attack directions.
Concretely, for Exp. 4 the objective is
$\mathcal{L}_{\text{Exp4}}=\mathcal{L}_{\text{mask}}+\mathcal{L}_{\text{logit}}+\mathcal{L}_{\text{cons}},$ (37)
where $\mathcal{L}_{\text{mask}}$ is the same masked-input cross-entropy term as in Exp. 3:
$\mathcal{L}_{\text{mask}}=\mathrm{CE}\!\big(f_{\theta}(X_{\mathrm{mask}}),y\big).$ (38)
The released script uses a shared logit-side stabilizer across ResNet-18, ViT-Tiny, and EfficientNet-B0, namely a centered-logit variance penalty. Let
$\bar{z}=z-\frac{1}{K}\sum_{j=1}^{K}z_{j},\qquad\mathrm{VarC}(z)=\frac{1}{K}\sum_{j=1}^{K}\bar{z}_{j}^{2}.$ (39)
Then
$\mathcal{L}_{\text{logit}}=\alpha_{\mathrm{logit}}\cdot\frac{1}{2}\Big(\mathrm{VarC}(f_{\theta}(X))+\mathrm{VarC}(f_{\theta}(X_{\mathrm{mask}}))\Big),$ (40)
where $\alpha_{\mathrm{logit}}=0$ at epoch 0, and $\alpha_{\mathrm{logit}}=\texttt{penalty\_weight}$ thereafter.
When the consistency branch is activated, the auxiliary view is a per-sample mean-baseline image:
$X_{\mathrm{aux}}=\bar{X}\mathbf{1},\qquad\bar{X}=\frac{1}{CHW}\sum_{c,h,w}X_{c,h,w}.$ (41)
The consistency term is
$\mathcal{L}_{\text{cons}}=0.8\,\mathrm{KL}\!\Big(\log\mathrm{softmax}\!\big(\tfrac{f_{\theta}(X_{\mathrm{mask}})}{T}\big)\,\big\|\,\mathrm{softmax}\!\big(\tfrac{f_{\theta}(X_{\mathrm{aux}})}{T}\big)\Big),$ (42)
with $T=1$. In the released implementation, this branch is enabled for epochs greater than 5.
We evaluate under the same $L_{\infty}$ setting as Exp. 3 ($\epsilon=8/255$) using PGD-20, APGD-CE, and APGD-DLR. The following observations can be made:
@@
    Lightweight stabilizers restore DLR robustness with practical efficiency. Adding shared logit regularization and shared auxiliary-view consistency on top of the same ConjMask procedure substantially improves APGD-DLR robustness where Exp. 3 remains weak. On ResNet-18, APGD-DLR increases from $1.28\%$ (Exp. 3) to $43.29\%$ (Exp. 4), close to TRADES-PGD’s $42.27\%$, while maintaining high clean accuracy ($87.64\%$) and requiring no adversarial training, hence lower practical cost. The resulting robustness profile is more balanced than Exp. 3, with non-trivial performance retained under PGD-20 and APGD-CE ($62.71\%$ and $36.92\%$, respectively).
@@
    The gain is architecture-dependent in outcome, reflecting CE–DLR direction mismatch with partial overlap. For ViT-Tiny and EfficientNet-B0, Exp. 4 trades part of the PGD/APGD-CE robustness of Exp. 3 for improved APGD-DLR robustness (ViT-Tiny: $13.96\%\to 15.91\%$; EfficientNet-B0: $0.20\%\to 28.36\%$), but the improvement remains less pronounced than on ResNet-18 and is accompanied by a noticeable drop in PGD-20/APGD-CE.
Overall, the shared LogitReg augmentation can deliver non-trivial robustness improvements with much less training time than adversarial training.
@@
We test Proposition 4 by testing whether low prompt-level conjugate correlation signals hallucination risk in a trained LLM. For each of 500 math problems, we compute the prefill-stage probe on the prompt embeddings: we obtain the embedding tensor $x$ and its input gradient $p=\nabla_{x}\mathcal{L}(x)$ from a single backward pass, and form $c_{\text{prompt}}$ (Eq. (34)), which is computed before any answer tokens are generated.
@@
We report AUROC with bootstrap resampling ($N_{\text{boot}}=2000$) for distinguishing strict-consensus hallucinations ($P_{\text{vote}}=1.0$) from strict-consensus clean responses ($P_{\text{vote}}=0.0$), and also provide a relaxed-consensus analysis ($P_{\text{vote}}\geq 0.8$) as a robustness check. Results are summarized in Table II. We can observe that:
@@
We use Perturbation-100, consisting of 100 problems with 5 semantically equivalent (paraphrased) prompt variants per problem. For each problem, we compute a prefill risk score for all five candidate prompts without decoding any answer tokens. The selection rule itself is therefore prefill-only: for each metric, the prompt with the lowest predicted risk is designated as the selected variant. For offline evaluation, we then generate deterministic answers for all five prompt variants and score them with the same judge panel as in Exp. 5. The resulting judged answers are used only to identify the judge-best set for each problem (allowing ties among highest-scoring variants) and to compute the Top-1 / Mean-Regret metrics reported below. As a sanity check on the reliability of the judge panel, we also report inter-judge agreement via Fleiss’ Kappa.
@@
Each candidate prompt is assigned one scalar risk value (higher $\rightarrow$ worse) computed before decoding: Risk-Cos (ours), Risk-Margin, Risk-Entropy, and Risk-Loss (NLL) (definitions follow Sec. IV-B). All reported selection decisions in this experiment use only these prefill scores.
@@
We focus on two selection-oriented metrics that directly reflect the practical utility of a prefill-only prompt filter:
  • Top-1 Hit Rate ($\uparrow$): the fraction of problems for which the selected prompt belongs to the judge-best set.
  • Mean Regret ($\downarrow$): the average quality gap between the selected prompt and the best judge score attained within the same five-prompt cluster.
In addition, we report Fleiss’ Kappa on the judges’ binary hallucination annotations as a sanity check on inter-judge consistency.
Prefill Risk Score Top-1 Hit Rate $\uparrow$ Mean Regret $\downarrow$
Risk-Cos (ours) 0.83 0.433
Risk-Margin 0.81 0.591
Risk-Loss (NLL) 0.78 0.615
Risk-Entropy 0.77 0.636
TABLE III: Prefill-only prompt selection results on Perturbation-100. For each five-prompt cluster, the prompt with the lowest predicted prefill risk is selected before decoding. Risk-Cos achieves the highest Top-1 Hit Rate and the lowest Mean Regret among all compared methods.